PHOTOELECTRON SPECTROSCOPY OF ClH_{2}^{−} AND ClD_{2}^{−}: A PROBE OF THE Cl + H_{2} VAN DER WAALS WELL AND SPIN-ORBIT EXCITED STATES

\underline{MICHAEL J. FERGUSON}, GIOVANNI MELONI, HARRY GOMEZ, and DANIEL M. NEUMARK, Department of Chemistry, University of California, Berkeley, CA 94720 and Chemical Sciences Division, Lawrence Berkeley National Laboratory, Berkeley, CA 94720.

The anion photoelectron spectra of ClH_{2}^{−} and ClD_{2}^{−} were measured at 299 nm, accessing the prereactive van der Waals well on the ground state Cl + H_{2} potential energy surface, as well as the low-lying spin-orbit excited states resulting from the interaction of Cl (^{2}P_{3/2}) and Cl^{∗} (^{2}P_{1/2}) with H_{2}. The photoelectron spectra are dominated by two relatively narrow peaks corresponding to the transitions to the neutral Cl·H_{2} and Cl^{∗}·H_{2} complexes both are shifted toward higher electron binding energy relative to atomic chlorine due to the solvent shift by the hydrogen. Another feature in the spectra is the relative broadening of ClD_{2}^{−} versus ClH_{2}^{−}. The broadening of the ClD_{2}^{−} peaks may have its origins in the excited state transitions. The broadening may also be in part due to the presence of two nuclear spin species, as opposed to one in ClH_{2}^{−}. Only ClH_{2}^{−} (ortho) was generated in this experiment, while ClD_{2}^{−} had both ortho and para complexes generated.